OpenStudy (anonymous):

Sketch one cycle of y = sin 2theta

OpenStudy (dan815):

so u get it...

OpenStudy (dan815):

i have to delete that.. there are people on this site that will give me warning and ban me for that picture

OpenStudy (dan815):

graph y= sin theta

OpenStudy (dan815):

how would multiplying the angle by 2 change it?

OpenStudy (ybarrap):

Fnd all multiples of \( \theta \) that make y zero. This marks the beginning and ends of one period. This turns out to be when \( 2\theta \) is zero or some multiple of \( 2\pi \). If you know the general characteristics of the sine wave, you know that it will either go up or down after one of these points. Generally something, like this, you'll just need to find out where y hits the x-axis: